OpenStudy (anonymous):

Solve the following for y: 2x + 2y = 8 y = −x + 4 y = x + 4 y = 3x y = −3x

OpenStudy (olyalhic):

2x +2y = 8 Add -2x to both sides. 2x+2y+−2x=8+−2x 2y=−2x+8 Divide both sides by 2. 2y2=−2x+82 y=−x+4 Answer y=−x+4

OpenStudy (anonymous):

ohhhhh thank you

OpenStudy (olyalhic):

@gabylovesyou100 your welcome